One measure of variation that we can look at is the Range when we have continuous variables. We will be using this measure to create a histogram. A histogram is used to show the frequencies of certain events or categories of data values in a set of data from one time period. Data are plotted in increasing or deceasing order based on the frequency count for each data categories.

Assignment I

Below is a data file containing the average number of minutes it took to code inpatient records each day in the health information management department. Data were collected for 30 days.

Average Time to Code Inpatient Records (minutes)
day min
1 30

2 31

3 30

4 27

5 32

6 33

7 31

8 29

9 28

10 34

11 29

12 30

13 27

14 31

15 28

16 30

17 32

18 29

19 31

20 30

21 28

22 26

23 28

24 30

25 33

26 30

27 31

28 29

29 32

30 31

Calculate the range in coding times (highest to lowest) for the time period and the average number of minutes for the time period.

Use your calculations to create a histogram of the data
Answer #1

To calculate range in coding times find the highest and lowest times

Highest time coded is 34 mins

Lowest time coded is 26 mins

Range= highest value - lowest value= 34-26= 8 mins

The average number of minutes for the time period= Total minutes /Total number of days= 26+27+27+28+28+28+28+29+29+29+29+30+30+30+30+30+30+30+31+31+31+31+31+31+32+32+32+33+33+34/30

= 900/30=30 mins